# Relations Among Different Inequality Measures in Complex Systems: From Kinetic Exchange to Earthquake Models

### v1

@@ 5.1 B-Model with Exchange Range
For each value of the exchange range parameter ($R$) used in the B-Model, we calculate all three indices and present the results in Table 1. Additionally, motivated from a recent work [asim_2025], we examine the ratios p/(2k1)p/(2k-1) and $p/g$ to understand how these derived quantities evolve with the $g$ and $k$, offering further insights into the internal consistency and comparative behavior of these inequality measures.
@@
Table 1: B-Model with Exchange Range: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-Model with different range $R$ computed from the Eqn (5). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different range ($R$).
Figure 2 provides a visual summary of how the inequality measures evolve as we vary the range parameter in the B-Model. In the upper-left panel, we can see the variation of $g$, $p$ and $k$ against $R$. In the upper-right panel, we illustrate the joint behaviors of the $g$-$k$ and $g$-$p$ pairs. As $R$ increases, $g$ and $k$ indices initially change in an almost perfectly linear manner with a slope $(3/8)$ as discussed in [bijin]. During this early phase, the values of $g$ and $k$ coincide at a point approximately $0.85$. Beyond this point, however, the relationship becomes distinctly nonlinear, indicating that the two measures respond differently to further increases in the range parameter. On the other hand as $R$ increases, $g$ and $p$ both increases together in similar manner and going to coincide at extreme inequality point $1$.
The lower-left panel of Figure 2 displays the variation of the ratio p/(2k1)p/(2k-1) with respect to $g$ and $k$. The pattern shows an initial increase of this ratio up to a certain threshold value; subsequently, as $g$ and $k$ continues to grow, the ratio begins to decline. Throughout the range, the ratio remains slightly above unity, reaching its maximum about a $3\%$ elevation near the point where $g$ and $k$ coincide (in this case, $g=k\sim 0.85$). This location has previously been identified as a precursor to criticality in several self-organized critical models [sand, manna, soumya]. The mild, non-monotonic rise toward this peak followed by a gradual decrease reflects a subtle shift in the balance between the $p$ and $k$ indices as the underlying wealth-exchange dynamics approach this near-critical region. A similar trend is observed in the lower-right panel, which shows how the ratio $p/g$ depends on $g$ and $k$ indices. In this case, the ratio increases sharply with increasing the indices $g$ and $k$, suggesting that $p$ grows at a faster rate relative to $g$.
@@ 5.2 B-Model with Savings
Table 2: B-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-model with different saving propensities $\lambda$ computed from the Eqn. (3.1.2). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different saving propensities $\lambda$.
@@ 5.3 C-Model with Savings
Table 3: C-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for C-model with different saving propensities $\lambda$ computed from the Eqn. (7). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for C-model with different saving propensities $\lambda$.
@@ 5.4 P-Model
We simulated income distributions governed by the Pareto law following Eqns. (8) and (9) and evaluated the associated inequality measures- namely, the Gini coefficient $g$, the Pietra index $p$, and the Kolkata index $k$ using Eqns. (1)–(4). The minimum income is fixed at unity, and all three indices exhibit a systematic and monotonic dependence on the Pareto shape parameter $\alpha$. As $\alpha$ approaches its lower bound of $1^{+}$, the distribution enters the regime of maximal permissible inequality while still maintaining a finite mean. Correspondingly, the indices converge toward their characteristic limiting values of approximately $0.85$.
A decrease in $\alpha$ enhances the heaviness of the power-law tail, implying an increasingly concentrated allocation of income among a diminishing fraction of individuals. This redistribution toward the upper tail is quantitatively captured by the pronounced rise in both $g$ and $k$ indices, each of which serves as a robust scalar measure of inequality. Their simultaneous escalation with decreasing $\alpha$ demonstrates that the broadening of the tail is directly reflected in the macroscopic inequality descriptors of the system.
@@
Table 4: P-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for P-model with different exponent $\alpha$, calculated using Eqn. (9), along with the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for P-model with different exponent $\alpha$.
@@ 5.5 CS-Model
The Kolkata–Gini relation obtained from the sequence of Cantor generations, initially exhibits small but systematic oscillations around an otherwise smooth and approximately monotonic curve. Here too, we calculate $g$, $k$ and $p$, along with the ratios p/(2k1)p/(2k-1) and $p/g$, to understand how inequality evolves with each fractal generation. The table 4 shows values at 5-generation intervals from generation 5 onward, with generation 4 included separately. .
@@
Table 5: CS-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for CS-model with different generations ($n$), where the overlap magnitude $Y_{n}(t)$ is calculated using Eqn. (10) along with the computed ratios p/(2k1)p/(2k-1) and $p/g$ (motivated by [asim_2025]).
@@ 5.6 BK-Model
Figure 7: BK-model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with number of blocks (NN) = $100$, $\alpha=1000$, $k_{c}/k_{p}=1$ and $\sigma=0.01$ for BK-Model. (a) Elastic energy released with respect to time ($t$); (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and also $g$ vs $p$-index curve; (c) pp/(2k1)(2k-1) vs. $g$ and $k$ and; (d) pp/$g$ vs. $g$ and $k$ plotted beyond the transient phase. Here the equation of motion of the blocks and the frictional force of them are given by Eqns. (11) and (12) respectively.
@@
In the bottom left panel we show p/(2k1)p/(2k-1), where the values drop below 1, ranging from $\approx 0.98$ to $\approx 1.08$İn the bottom right panel, we plot $p/g$ after the initial transient period (first 100 steps) has passed and a steady state is reached, since the system starts. In the Burridge-Knopoff model, the initial evolution is dominated by stress-loading transients arising from the imposed initial conditions. During this phase, event statistics and associated inequality measures such as the Gini and Kolkata indices exhibit strong time dependence and are not representative of the intrinsic dynamics. In the bottom right panel, the ratio $p/g$ ranges from $\approx 0.69$ and to $\approx 0.82$
6 Summery and Concluding remarks
@@
In this study, we have performed a comparative numerical investigation of inequality measures (namely of Gini [gini], Pietra [pietra] and Kolkata [k-index1, k-index3] indices) across a broad spectrum of model systems: they range from two kinetic wealth-exchange models (namely the Banerjee Model [B_Model, B_Model1] and Chakraborti model [C_Model], more popularly known as Yard-Sale model [YS, YS1]), to synthetic Pareto model [Pareto, Pareto1] and two earthquake-generating models (namely the Chakrabatrti-Stinchcombe model [CS_Model, CS1, pbhattacharya] and the Burridge-Knopoff model [langer_1994, BK_Model]). Using the Lorenz curves (see Fig. 1), generated numerically for each of these model systems for different parameter values (using Eqns. (5),(3.1.2) for B-Model, Eqn. (7) for C-Model, Eqn. (9) for P-Model, Eqn. (10) for CS-Model and Eqns. (11),(12) for the BK-Model), we estimate the different inequality indices, namely the Gini index ($g$), Pietra index ($p$), and Kolkata index ($k$). The results of our numerical analysis for the above mentioned inequality indices are given and shown respectively in Tables 1-5 and Figs. 2-7. It may be noted that while switching from the econophysical wealth exchange models (B and C models) which have got dominantly exponential (to Gamma-like) distributions and the two earthquake simulating models (CS and BK models) which have got (Guttenberg-Richter like) dominantly power law distributions, we considered and compared also the results for the perfect power-law Pareto distribution as well.
@@
We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models, the two-fractal model and the Pareto distributions) remain a little above unity, which was predicted to be exactly equal to unity [asim_2025]. The value of this quantity for the two socio-economic (B and C models) and the earthquake simulating CS model is seen to deviate a little above unity, bounded by a maximum of $4\%$ near $g=k=0.86\pm 0.02$ which was identified earlier [manna, soumya] to be the precursor point of criticality in several self-organized critical models. For the Pareto distribution, the ratio p/(2k1)p/(2k-1) deviates slightly above unity, reaching a maximum of about $7\%$ (see Fig. 5(c)). In the BK-Model, however, the value of p/(2k1)p/(2k-1) even drops below unity in certain regimes (see Fig. 7(c)). The predicted values (assuming a simple polynomial form for the Lorenz function [asim_2025]) of $p/g=3/4$ seems to deviate quite a bit (in the range $0.69$ to about $1.0$; see Tables 1-5), while the existence of the relation $k=1/2+(3/8)g$ [asim_2025] seems to be valid for all the models considered here in the region of lower $g$ values (see Figs. 2(d) - 7(d)).

### v2

@@ 5.1 B-Model with Exchange Range
For each value of the exchange range parameter ($R$) used in the B-Model, we calculate all three indices and present the results in Table 1. Additionally, motivated from a recent work [19], we examine the ratios p/(2k1)p/(2k-1) and $p/g$ to understand how these derived quantities evolve with the $g$ and $k$, offering further insights into the internal consistency and comparative behavior of these inequality measures.
@@
Table 1: B-Model with Exchange Range: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-Model with different range $R$ computed from the Eqn (5). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for B-model with different range ($R$).
Figure 2 provides a visual summary of how the inequality measures evolve as we vary the range parameter in the B-Model. In the upper-left panel, we can see the variation of $g$, $p$ and $k$ against $R$. In the upper-right panel, we illustrate the joint behaviors of the $g$-$k$ and $g$-$p$ pairs. As $R$ increases, $g$ and $k$ indices initially change in an almost perfectly linear manner with a slope $(3/8)$ as discussed in [22]. During this early phase, the values of $g$ and $k$ coincide at a point approximately $0.85$. Beyond this point, however, the relationship becomes distinctly nonlinear, indicating that the two measures respond differently to further increases in the range parameter. On the other hand as $R$ increases, $g$ and $p$ both increases together in similar manner and going to coincide at extreme inequality point $1$.
The lower-left panel of Figure 2 displays the variation of the ratio p/(2k1)p/(2k-1) with respect to $g$ and $k$. The pattern shows an initial increase of this ratio up to a certain threshold value; subsequently, as $g$ and $k$ continues to grow, the ratio begins to decline. Throughout the range, the ratio remains slightly above unity, reaching its maximum about a $3\%$ elevation near the point where $g$ and $k$ coincide (in this case, $g=k\sim 0.85$). This location has previously been identified as a precursor to criticality in several self-organized critical models [2, 25, 15]. The mild, non-monotonic rise toward this peak followed by a gradual decrease reflects a subtle shift in the balance between the $p$ and $k$ indices as the underlying wealth-exchange dynamics approach this near-critical region. A similar trend is observed in the lower-right panel, which shows how the ratio $p/g$ depends on $g$ and $k$ indices. In this case, the ratio increases sharply with increasing the indices $g$ and $k$, suggesting that $p$ grows at a faster rate relative to $g$.
@@ 5.2 B-Model with Savings
Table 2: B-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-model with different saving propensities $\lambda$ computed from the Eqn. (3.1.2). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for B-model with different saving propensities $\lambda$.
@@ 5.3 C-Model with Savings
Table 3: C-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for C-model with different saving propensities $\lambda$ computed from the Eqn. (7). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for C-model with different saving propensities $\lambda$.
@@ 5.4 P-Model
We simulated income distributions governed by the Pareto law following Eqns. (8) and (9) and evaluated the associated inequality measures- namely, the Gini coefficient $g$, the Pietra index $p$, and the Kolkata index $k$ using Eqns. (1)–(4). The minimum income is fixed at unity, and all three indices exhibit a systematic and monotonic dependence on the Pareto shape parameter $\alpha$. As $\alpha$ approaches its lower bound of $1^{+}$, the distribution enters the regime of maximal permissible inequality while still maintaining a finite mean. Correspondingly, the indices converge toward their characteristic limiting values. A decrease in $\alpha$ enhances the heaviness of the power-law tail, implying an increasingly concentrated allocation of income among a diminishing fraction of individuals. This redistribution toward the upper tail is quantitatively captured by the pronounced rise in both $g$ and $k$ indices, each of which serves as a robust scalar measure of inequality. Their simultaneous escalation with decreasing $\alpha$ demonstrates that the broadening of the tail is directly reflected in the macroscopic inequality descriptors of the system.
@@
Table 4: P-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for P-model with different exponent $\alpha$, calculated using Eqn. (9), along with the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for P-model with different exponent $\alpha$.
@@ 5.5 CS-Model
The Kolkata–Gini relation obtained from the sequence of Cantor generations, initially exhibits small but systematic oscillations around an otherwise smooth and approximately monotonic curve. Here too, we calculate $g$, $k$ and $p$, along with the ratios p/(2k1)p/(2k-1) and $p/g$, to understand how inequality evolves with each fractal generation. The table 4 shows values at 5-generation intervals from generation 5 onward, with generations 4 and 199 included separately.
@@
Table 5: CS-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for CS-model with different generations ($n$), where the overlap magnitude $Y_{n}(t)$ is calculated using Eqn. (10) along with the computed ratios p/(2k1)p/(2k-1) and $p/g$ (motivated by [19]).
@@ 5.6 BK-Model
Figure 7: BK-model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with number of blocks (NN) = $100$, $\delta=1000$, $k_{c}/k_{p}=1$ and $\sigma=0.01$ for BK-Model. (a) Elastic energy released with respect to time ($t$); (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and also $g$ vs $p$-index curve; (c) pp/(2k1)(2k-1) vs. $g$ and $k$ and; (d) pp/$g$ vs. $g$ and $k$ plotted beyond the transient phase. Here the equation of motion of the blocks and the frictional force of them are given by Eqns. (11) and (12) respectively.
@@
In the bottom left panel we show p/(2k1)p/(2k-1), where the values drop below 1, ranging from $\approx 0.98$ to $\approx 1.08$ . In the bottom right panel, we plot $p/g$ after the initial transient period (first 100 steps) has passed and a steady state is reached, since the system starts. In the Burridge-Knopoff model, the initial evolution is dominated by stress-loading transients arising from the imposed initial conditions. During this phase, event statistics and associated inequality measures such as the Gini and Kolkata indices exhibit strong time dependence and are not representative of the intrinsic dynamics. In the bottom right panel, the ratio $p/g$ ranges from $\approx 0.69$ and to $\approx 0.82$
6 Summary and concluding remarks
In this study, we have performed a comparative numerical investigation of inequality measures (namely of Gini [21], Pietra [29] and Kolkata [14, 3] indices) across a broad spectrum of model systems: they range from two kinetic wealth-exchange models (namely the Banerjee Model [5, 18] and
@@ 6 Summary and concluding remarks
Chakraborti model [12], more popularly known as Yard-Sale model [8, 7]), to synthetic Pareto model [28, 26] and two earthquake-generating models (namely the Chakrabatrti Stinchcombe model [10, 23, 6] and the Burridge Knopoff model [9, 27]). Using the Lorenz curves (see Fig. 1), generated numerically for each of these model systems for different parameter values (using Eqns. (5),(3.1.2) for B-Model, Eqn. (7) for C-Model, Eqn. (9) for P-Model, Eqn. (10) for CS-Model and Eqns. (11),(12) for the BK-Model), we estimate the different inequality indices, namely the Gini index ($g$), Pietra index ($p$), and Kolkata index ($k$). The results of our numerical analysis for the above mentioned inequality indices are given and shown respectively in Tables 1-5 and Figs. 2-7. It may be noted that while switching from the econophysical wealth exchange models (B and C models) which have got dominantly exponential (to Gamma-like) distributions and the two earthquake simulating models (CS and BK models) which have got (Gutenberg-Richter like) dominantly power law distributions, we considered and compared also the results for the perfect power-law Pareto distribution as well.
@@
We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models, the two-fractal model and the Pareto distributions) remain a little above unity, which was predicted to be exactly equal to unity [19]. The value of this quantity for the two socio-economic (B and C models) and the earthquake simulating CS model is seen to deviate a little above unity, bounded by a maximum of $4\%$ near $g=k=0.86\pm 0.02$ which was identified earlier [25, 15] to be the precursor point of criticality in several self-organized critical models. For the Pareto distribution, the ratio p/(2k1)p/(2k-1) deviates slightly above unity, reaching a maximum of about $7\%$ (see Fig. 5(c)). In the BK-Model, however, the value of p/(2k1)p/(2k-1) reaches a maximum of about $8\%$ above unity (see Fig. 7(c)). The predicted values (assuming a simple polynomial form for the Lorenz function [19]) of $p/g=3/4$ seems to deviate quite a bit (in the range $0.69$ to about $1.0$; see Tables 1-5), while the existence of the relation $k=1/2+(3/8)g$ [19] seems to be valid for all the models considered here in the region of lower $g$ values (see Figs. 2(d) - 7(d)).